LASER-INDUCED PHOSPHORESCENCE OF SO_{2} IN SOLID NEON : DIRECT OBSERVATION OF THE b 3A_{2} STATE

\underline{YUAN-PERN LEE}, Department of Chemistry, National Tsing Hua University, Hsinchu, Taiwan 30043; A. J. MERER, Department of Chemistry, The University of British Columbia, 2036 Main Mall, Vancouver, B. C., Canada V67 1Z1.

Dispersed phosphorescence spectra of the transition a ^{3}B_{1} - X ^{1}A_{1} of ^{18}O-labeled SO_{2} in solid Ne were recorded in the spectral region 388-478 nm during laser excitation. As zero-phonon lines are well resolved, we were able to identify all three vibrational modes of the X ^{1}A_{1} state, and to monitor emission from a specific isotopomer (in particular ^{16}OS^{18}O) to obtain its corresponding excitation spectra. The vibronic patterns of excitation spectra of S^{16}O_{2} and S^{18}O_{2} in the spectral range 351-388 nm are similar to those observed in the gaseous phase, with severely perturbed lines above a ^{3}B_{1} (110) at 27005 cm^{−1}. New lines in two groups were observed for ^{16}OS^{18}O; they are identified as lines associated with the ν_{3} of a ^{3}B_{1} and of b ^{3}A_{2}. These observations enable us to report ν_{3} = 922 cm^{−1} for a ^{3}B_{1}, ν_{2} = 315 and ν_{1} = 771 cm^{−1} for b ^{3}A_{2}, and a false origin at 26653 cm^{−1} for the b ^{3}A_{2} (001) - X ^{1}A_{1} (000) transition of ^{16}OS^{18}O.